# DreamOmni: Unified Image Generation and Editing

### v1

@@ 1 Introudction
After unified training with T2I data and various synthetic datasets, our DreamOmni showcases competitive T2I generation capabilities and delivers outstanding performance across a range of editing tasks.
@@ 2 Related Work
Text-to-image Diffusion Model. Diffusion models [52, 23, 14, 24, 1, 21, 5, 53, 35] have emerged as highly effective methods for image generation, outperforming previous generative models like GANs [20] and VAEs [26]. The latent diffusion model [44], also known as Stable Diffusion, enhances the diffusion process within the image latent space and utilizes cross-attention to guide image generation through text input. This approach enhances the usability of the image generation model, driving its widespread adoption. Subsequently, more advanced text-to-image (T2I) models were developed, such as Imagen [47], DALLE [43], and others [6, 3, 16, 39, 13]. Recently, the DIT frameworks [38, 4] has become increasingly popular as more models [15, 10, 31] transition from the Unet framework [45], thanks to its straightforward structure and outstanding compatibility with multimodal inputs. However, the current T2I model fails to account for the design of downstream editing tasks, necessitating numerous specialized adaptations. For instance, Instructp2p [8] requires an expansion of input channels, while ControlNet [62] and IP-adapter [60] rely on additional plugins and various methods for injecting reference image information. These fragmented designs are inconsistent with the trend toward unified models, complicating deployment and impeding the joint training of T2I and editing tasks. In this paper, we will examine and analyze the existing T2I model frameworks and propose a more efficient and powerful unified generation and editing model.
@@ 3.1 Framework
In this part, we aim to design a unified and powerful image generation and editing framework. Currently, different editing models often have distinct structure designs. For example, IP-adapter [60] and BLIP-Diffusion [28] inject information through cross-attention to maintain subjects. In contrast, InstructP2P [8] achieves editing consistency by adding different numbers of input channels for models. These structures are tailored for specific tasks and lack generalizability. To this end, as shown in Fig. 2 (a), we concatenate VLM features with noisy latent and input them into a DIT Block for joint multi-head self-attention operations. After that, the VLM features and noisy latent are processed by the FeedForward modules. This allows the model to autonomously learn any level features (from overall consistency to subject consistency) for editing and generation. Notably, for FeedForward modules, we separate the VLM features and noisy latent, passing them through two distinct FeedForward modules with the same network structure. Additionally, instead of using CLIP [41] or T5 [42] as the text encoder, we introduce a Vision-Language Model (VLM) that enables joint understanding and encoding of both image and text prompts.
@@ 3.2 Synthetic Data
Overall, our synthetic collage data pipeline is both efficient and precise, enabling the generation of billions of diverse edited images with ease. It is highly suitable for large-scale pretraining and fine-tuning of DreamOmni.

### v2

@@ 1 Introudction
After unified training with T2I data and various synthetic datasets, our DreamOmni showcases competitive performance on T2I generation and various editing tasks.
@@ 2 Related Work
Text-to-image Diffusion Model. Diffusion models [52, 23, 14, 24, 1, 21, 5, 53, 35] have emerged as highly effective methods for image generation, outperforming previous generative models like GANs [20] and VAEs [26]. The latent diffusion model [44], also known as Stable Diffusion, enhances the diffusion process within the image latent space and utilizes cross-attention to guide image generation through text input. This approach enhances the usability of the image generation model, driving its widespread adoption. Subsequently, more advanced text-to-image (T2I) models were developed, such as Imagen [47], DALLE [43], and others [6, 3, 16, 39, 13]. Recently, the DIT frameworks [38, 4] has become increasingly popular as more models [15, 10, 31] transition from the Unet framework [45], thanks to its straightforward structure and outstanding compatibility with multimodal inputs. However, the current T2I model fails to account for the design of downstream editing tasks, necessitating numerous specialized adaptations. For instance, Instructp2p [8] requires an expansion of input channels, while ControlNet [62] and IP-adapter [60] rely on additional plugins and various methods for injecting reference image information. These fragmented designs are inconsistent with the trend toward unified models, complicating deployment and impeding the joint training of T2I and editing tasks. In this paper, we will analyze the existing T2I frameworks and propose a more efficient and powerful unified generation and editing model.
@@ 3.1 Framework
In this part, we aim to design a unified and powerful image generation and editing framework. Currently, different editing models often have distinct structure designs. For example, IP-adapter [60] and BLIP-Diffusion [28] inject information through cross-attention to maintain subjects. In contrast, InstructP2P [8] achieves editing consistency by adding different numbers of input channels for models. These structures are tailored for specific tasks and lack generalizability. To this end, as shown in Fig. 2 (a), we concatenate VLM features with noisy latent and input them into a DIT Block for joint multi-head self-attention operations. After that, the VLM features and noisy latent are processed by the FeedForward modules. This allows the model to autonomously learn any level features (from overall consistency to subject consistency) for editing and generation. Notably, for FeedForward modules, we separate the VLM features and noisy latent, passing them through two distinct FeedForward modules with the same network structure. Additionally, instead of using CLIP [41] or T5 [42] as the text encoder, we introduce a Vision-Language Model (VLM) that enables joint understanding and encoding of both image and text prompts. For tasks that require high consistency, such as instruction-based editing and drag editing, we input the VAE encoding of the source image into the DIT model to ensure strong consistency between the non-edited regions of the output image and the source image.
@@ 3.2 Synthetic Data
Our synthetic collage data pipeline is efficient and precise, enabling the generation of billions of diverse images for large-scale pretraining and fine-tuning of DreamOmni.
3.3 Model Training